OpenStudy (anonymous):

Solve 4sin(x)+3cos(x) = 2 for the first 2 positive solutions

OpenStudy (anonymous):

Refer to the attached solution using Mathematica 9.

OpenStudy (dumbcow):

Divide by cos \[4 \tan x + 3 = 2 \sec x\] square both sides \[16 \tan^2 x + 24 \tan x + 9 = 4 \sec^2 x\] use identity: sec^2 = 1+tan^2 \[16 \tan^2 x + 24 \tan x + 9 = 4 + 4 \tan^2 x\] \[12 \tan^2 x + 24 \tan x + 5 = 0\] \[\tan x = \frac{-24 \pm \sqrt{24^2 - 4(12)(5)}}{2(12)}\] simplified \[\tan x = -1 \pm \frac{\sqrt{21}}{6}\] \[x = \pi n + \tan^{-1} ( -1 \pm \frac{\sqrt{21}}{6})\] From here use a calculator to estimate the first 2 positive solutions

OpenStudy (dumbcow):

sorry I forgot to add you need to also check for extraneous solutions the above solution gives the following positive solutions \[x \approx 2.08, 2.91, 5.22, 6.05\] two are extraneous, dont work if you plug into original equation real solutions: \[x \approx 2.08, 6.05\]

OpenStudy (anonymous):

@dumbcow Nice work. Mathematica allows someone like myself, 80 years old this August, to particpate on this site. I'm not sure what the derivative of x^2 is. At this time I cannot remember any of the derivative rules. Somewhere in that 1.5 Gb installation file, Mathematica has access to all of them.